OpenStudy (anonymous):

Find F^-1(x) f(x)=ln(x+3)

OpenStudy (anonymous):

is this because ln = e that we have e^(x+3)

OpenStudy (anonymous):

or is it because the inverse of ln is e

OpenStudy (anonymous):

im kinda confused on that one

OpenStudy (kropot72):

\[f(x)=y=\ln (x+3)\] Interchanging x and y to find the inverse: \[x=\ln (y+3)\] Therefore \[e ^{x}=y+3\] \[f ^{-1}(x)=e ^{x}-3\]

OpenStudy (anonymous):

Ok I understand now ty

OpenStudy (kropot72):

You're welcome :)